# Hartman Effect from a Geometrodynamic Extension of Bohmian Mechanics

### v2

@@ 4 Construction of the General Solution to the Problem
$\hat{g}_{\mu\nu}=\begin{cases}g_{\mu\nu}&\mbox{ for}x\leq-a/2\\ \mathcal{G}_{\mu\nu}&\mbox{for}-a/2<x<a/2\\ g_{\mu\nu}&\mbox{for}x\geq a/2\end{cases},$ (37)
written in terms of metrics $g_{\mu\nu}$ and $\mathcal{G}_{\mu\nu}$ which by construction reduces to a Minkowski-type metric when $Q\rightarrow mc^{2}/3=1/3$ (since in this article $c=1$, $\hbar=1$ and $m=1$) and $f(r_{s})\rightarrow 0$, which ultimately describes the entire spacetime through which particles propagate.
@@
$dS=-2dr_{s}v_{s}f\left\{\frac{1}{1-v^{2}_{s}f^{2}}\right\},$ (38)
@@
is obtained, whose solution will be presented later, allowing us to obtain the expression for the quantum potential in region $II$. Thus, by integrating Eq. (38), as shown in the part $I$ of the Appendix, the phase function is obtained:
$\begin{split}S_{II}(x,x_{s}(t))&=\frac{\alpha_{0}\beta_{0}}{2\sigma\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}}\Biggl{\{}\frac{\beta_{1}-\mu_{0}}{\sqrt{\mu_{0}}}\tan^{-1}{\Biggl{[}\frac{1}{\sqrt{\mu_{0}}}\tan{\frac{x^{\prime}}{2}}\Biggr{]}}-\frac{\beta_{1}-\mu_{1}}{\sqrt{\mu_{1}}}\tan^{-1}{\Biggl{[}\frac{1}{\sqrt{\mu_{1}}}\tan{\frac{x^{\prime}}{2}}\Biggr{]}}\Biggr{\}}.\end{split}$ (40)
Above $\alpha_{0}=-[v_{s}\tanh{(2\sigma R)}]/[2\tanh{(\sigma R)}]$, $\alpha_{1}=\mbox{sech}{(2\sigma R)}$, $\beta_{0}=[2(\alpha_{1}-1)]/[(\alpha_{1}-1)^{2}-\alpha^{2}_{0}]$, $\beta_{1}=(\alpha_{1}+1)/(\alpha_{1}-1)$, $\beta_{2}=[(\alpha_{1}+1)^{2}-\alpha^{2}_{0}]/[(\alpha_{1}-1)^{2}-\alpha^{2}_{0}]$, $\beta_{3}=[\alpha_{1}^{2}-1+\alpha^{2}_{0}]/[(\alpha_{1}-1)^{2}-\alpha^{2}_{0}]$, $\mu_{0}=\beta_{3}-\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}$, $\mu_{1}=\beta_{3}+\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}$ and
@@
$\Psi_{II}(x,t)=\sum_{l=1}^{2}C_{II}^{l}\sqrt{\rho_{II}^{l}}\exp{\{iS_{II}^{l}(x,t)\}}.$ (43)
However, it is important to note that physically acceptable solutions for Eq. (43), leading to a consistent phase function (40), are those where $\alpha_{1}\neq 0$ or $\alpha_{1}\neq 1$. Then, imposing the continuity conditions of the solution and its derivatives, the expressions obtained from Eqs. (29) and (43) must be substituted in
$\begin{split}\Psi_{k}(x,t)\biggl{|}_{(x_{k},t_{k})}=\Psi_{k+I}(x,t)\biggl{|}_{(x_{k},t_{k})}\;\;\;\mbox{and}\;\;\;[\partial_{x}\Psi_{k}(x,t)]\biggl{|}_{(x_{k},t_{k})}=[\partial_{x}\Psi_{k+I}(x,t)]\biggl{|}_{(x_{k},t_{k})},\end{split}$ (44)
with $x_{k}\in\{-a/2,a/2$, $t_{k}\in\{t_{0},t_{1}\}$ and $k=I,II,III$. Thus one obtains the system of equations:
$\begin{split}&\mathcal{C}_{I}^{1}\zeta_{1}+\mathcal{C}_{I}^{2}\zeta_{2}e^{i\Delta Et_{0}}=\zeta_{3}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{0}}],\\ &\mathcal{C}_{I}^{1}\zeta_{1}^{\prime}+\mathcal{C}_{I}^{2}\zeta_{2}^{\prime}e^{i\Delta Et_{0}}=\zeta_{3}^{\prime}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{0}}],\\ &\mathcal{C}_{III}^{1}\zeta_{5}+\mathcal{C}_{III}^{2}\zeta_{6}e^{i\Delta Et_{1}}=\zeta_{4}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{1}}],\\ &\mathcal{C}_{III}^{1}\zeta_{5}^{\prime}+\mathcal{C}_{III}^{2}\zeta_{6}^{\prime}e^{i\Delta Et_{1}}=\zeta_{4}^{\prime}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{1}}].\end{split}$ (45)
@@
$\begin{split}&\mathcal{C}_{I}^{1}=\varpi_{4}e^{-i\Delta Et_{0}}[(e^{i\Delta Et_{0}}-1)\mathcal{C}_{II}^{1}+1],\\ &\mathcal{C}_{I}^{2}=\varpi_{3}[(e^{i\Delta Et_{0}}-1)\mathcal{C}_{II}^{1}+1],\\ &\mathcal{C}_{II}^{1}=-\frac{e^{i\Delta Et_{1}}}{e^{i\Delta Et_{1}}-1}\left\{\varpi_{2}\mathcal{C}_{III}^{1}-1\right\},\\ &\mathcal{C}_{II}^{2}=\frac{1}{e^{i\Delta Et_{1}}-1}\left\{\varpi_{2}e^{i\Delta Et_{1}}\mathcal{C}_{III}^{1}-1\right\},\\ &\mathcal{C}_{III}^{2}=\varpi_{1}e^{i\Delta Et_{1}}\mathcal{C}_{III}^{1}.\end{split}$ (46)
@@
$\begin{split}&\zeta_{l}=\sqrt{\rho_{I}^{l}(x,t)}\exp{\{iS_{I}^{l}(x,t)\}}\biggl{|}_{(-\frac{a}{2},t_{0})},\\ &\zeta_{l}^{\prime}=\frac{k^{l}}{2\sqrt{\rho_{I}^{l}(x,t)}}[(\mathcal{B}^{2}-\mathcal{A}^{2})\sin{\{2p_{\mu}x^{\mu}\}}+i2\mathcal{A}\mathcal{B}]\exp{\{iS_{I}^{l}(x,t)\}}\biggl{|}_{(-\frac{a}{2},t_{0})},\\ &\zeta_{l+2}=\exp{\{i[S_{II}(x,x_{s}(t))-V_{0}t]\}}\biggl{|}_{((-1)^{l+2}\frac{a}{2},t_{0})},\\ &\zeta_{l+2}^{\prime}=\frac{i\alpha_{0}\beta_{0}\sec^{2}{(\frac{x^{\prime}}{2})}\tan{(2\sigma r_{s})}}{2\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}}\frac{\exp{\{i[S_{II}(x,x_{s}(t))-V_{0}t]\}}}{\sqrt{\cosh^{2}{(2\sigma r_{2})}-1}}\biggl{[}\frac{\beta_{1}-\mu_{1}}{\mu_{1}+\tan^{2}{(\frac{x^{\prime}}{2})}}-\frac{\beta_{1}-\mu_{0}}{\mu_{0}+\tan^{2}{(\frac{x^{\prime}}{2})}}\biggr{]}\biggl{|}_{((-1)^{l+2}\frac{a}{2},t_{0})},\\ &\zeta_{l+4}=\exp{\{iS_{III}^{l}(x,t)\}}\biggl{|}_{(\frac{a}{2},t_{0})},\\ &\zeta_{l+4}^{\prime}=ik^{l}\exp{\{iS_{III}^{l}(x,t)\}}\biggl{|}_{(\frac{a}{2},t_{0})}.\;\end{split}$ (47)
In the expressions above, values for $l=1,2$ are considered, along with the energy variation $\Delta E=E^{2}-E^{1}$. Additionally, the coefficients were rewritten as $C_{j}^{l}=\mathcal{C}_{j}^{l}$ if $j=I$ and $C_{j}^{l}\sqrt{\rho_{j}^{l}}=\mathcal{C}_{j}^{l}$ for $j=II,III$, again for all $l=1,2$.
@@
$\begin{split}\Psi_{I}(x,t)&=\mathcal{C}_{I}^{1}\sqrt{\rho_{I}^{1}}\exp{\{iS_{I}^{1}(x,t)\}}+\mathcal{C}_{I}^{2}\sqrt{\rho_{I}^{2}}\exp{\{iS_{I}^{2}(x,t)\}},\\ \Psi_{II}(x,t)&=[\mathcal{C}_{II}^{1}\exp{\{iE^{1}t\}}+\mathcal{C}_{II}^{2}\exp{\{iE^{2}t\}}]\exp{\{i[S_{II}(x,x_{s}(t))-V_{0}t]\}},\\ \Psi_{III}(x,t)&=\mathcal{C}_{III}^{1}\exp{\{iS_{III}^{1}(x,t)\}}+\mathcal{C}_{III}^{2}\exp{\{iS_{III}^{2}(x,t)\}},\end{split}$ (48)
@@ 5 General Expression for the Quantum Potential
For this purpose, it is necessary to recall that the metric describing spacetime in regions $I$ and $III$ implies that the quantum potential in these regions must be of the form $Q_{j}=\pm mc^{2}\{g_{00}\}^{3/2}/3=\pm\{g_{00}\}^{3/2}/3$, where again $j=I$ as $x\leq-a/2$ and $j=II$ as $x\geq a/2$. However, as mentioned in the previous section, the quantum potential in region II is determined by integrating (39), resulting in the expression
$Q_{II}(x,t)$ =$=\Biggl{\{}\ln{\Biggl{|}\frac{1+v_{s}f}{\sqrt{1-v^{2}_{s}f^{2}}}\Biggr{|}}+\frac{v^{2}_{s}f^{2}}{2}\Biggr{\}}.$ (50)
@@ 6 Momentum, Position, and Tunneling Time
In the previous sections, along with the theoretical foundations and the general solution, we derived the mathematical expressions for the quantum potential in each of the regions of interest in our problem. Even though the wave function and the quantum potential are fundamental quantities within the framework of Bohmian mechanics [33, 32, 42, 40], it is essential to determine the particle’s momentum and position. Therefore, this section is dedicated to calculating the mathematical expressions for these quantities, ultimately allowing us to determine the quantum tunneling time [1, 10, 29, 6, 7, 12, 13, 14].
In this regard, obtaining an expression for the linear momentum for the region $I$ turns out to be a difficult task due to the form of the general solution. Nevertheless, by rewriting $\sqrt{\rho_{I}^{l}}$ in the form
$\begin{split}\sqrt{\rho_{I}^{l}}=\sqrt{\mathcal{A}^{2}\cos^{2}{\{p_{\mu}x^{\mu}/\hbar\}}+\mathcal{B}^{2}\sin^{2}{\{p_{\mu}x^{\mu}/\hbar\}}}=\sqrt{2AB+[A^{2}+B^{2}]\cos{\{2p_{\mu}x^{\mu}\}}}\end{split}$ (51)
@@
$\begin{split}\sqrt{\rho_{I}^{l}}&=\sqrt{2AB+[A^{2}+B^{2}]\cos{\{2p_{\mu}x^{\mu}\}}}=\sqrt{(A+B)^{2}}=\mathcal{A}.\end{split}$ (52)
On the other hand, the phase function $S_{I}^{l}$ for low energies reduces to $S_{I}^{l}(x,t)=\tan^{-1}{\{\mathcal{B}\tan{\{p_{\mu}x^{\mu}\}}/\mathcal{A}\}}\approx\mathcal{B}\{k^{l}x-E^{l}t\}/\mathcal{A}$. Combining this simplified expression with equation (29) gives the phase function $S_{I}(x,t)$. Then, to determine the linear momentum in region $I$, we choose the metric in this region to be of Minkowski type, that is, $Q_{I}=mc^{2}/3$, which implies that Eq. (20) reduces to Eq. (7) when particles incident on the potential barrier have non-relativistic velocities ($v\ll 1$, taking natural units $\hbar=c=1$) which is precisely our case of study. Then, substituting $S_{I}(x,t)$ in Eq. (7), we obtain
$\begin{split}P_{I}(x,t)&=\frac{\mathcal{B}[(\mathcal{C}_{I}^{1})^{2}k^{1}+(\mathcal{C}_{I}^{2})^{2}k^{2}+\mathcal{C}_{I}^{1}\mathcal{C}_{I}^{2}(k^{1}+k^{2})\cos{\Xi_{I}}]}{\mathcal{A}[(\mathcal{C}_{I}^{1})^{2}+(\mathcal{C}_{I}^{2})^{2}+2\mathcal{C}_{I}^{1}\mathcal{C}_{I}^{2}\cos{\Xi_{I}}]}.\end{split}$ (53)
@@
$\begin{split}P_{III}(x,t)&=\frac{(\mathcal{C}_{III}^{1})^{2}k^{1}+(\mathcal{C}_{III}^{2})^{2}k^{2}+\mathcal{C}_{III}^{1}\mathcal{C}_{III}^{2}(k^{1}+k^{2})\cos{\Xi_{III}}}{(\mathcal{C}_{III}^{1})^{2}+(\mathcal{C}_{III}^{2})^{2}+2\mathcal{C}_{III}^{1}\mathcal{C}_{III}^{2}\cos{\Xi_{III}}}.\end{split}$ (54)
@@
$\begin{split}&P_{II}(x,t)=\frac{\alpha_{0}\beta_{0}\mbox{sech}(2\sigma r_{s})}{2\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}}\frac{1-\alpha_{0}^{2}+2\alpha_{1}\cosh{(2\sigma r_{s})}+\alpha_{1}^{2}\cosh^{2}{(2\sigma r_{s})}}{1+2\alpha_{1}\cosh{(2\sigma r_{s})}+\alpha_{1}^{2}\cosh^{2}{(2\sigma r_{s})}}\\ &\hskip 99.58464pt\times\Biggl{\{}\frac{\beta_{1}-\mu_{0}}{\mu_{0}\cos^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}+\sin^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}}-\frac{\beta_{1}-\mu_{1}}{\mu_{1}\cos^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}+\sin^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}}\Biggr{\}}.\end{split}$ (55)

### v3

@@ 4 Construction of the General Solution to the Problem
$\hat{g}_{\mu\nu}=\left\{\begin{array}[]{ccc}g_{\mu\nu}&\mbox{for}&x\leq-a/2,\\ \mathcal{G}_{\mu\nu}&\mbox{for}&-a/2<x<a/2,\\ g_{\mu\nu}&\mbox{for}&x\geq a/2,\end{array}\right$.
which is clearly continuous at the boundaries, since, at the entry points $\mathcal{P}_{0}(-a/2,t_{0})$ and exit points $\mathcal{P}_{1}(a/2,t_{1})$ the perturbation velocity must be null ($v_{s}$), therefore $\mathcal{G}_{\mu\nu}=\mbox{diag}=(-1,1,1,1)$. Furthermore, by construction we know that $g_{\mu\nu}$ reduces to a Minkowski-type metric when $Q\rightarrow mc^{2}/3=1/3$ (recalling that in this article $c=1$, $\hbar=1$ and $m=1$) for the description to be consistent. Thus, $g_{\mu\nu}=\mathcal{G}_{\mu\nu}=\mbox{diag}=(-1,1,1,1)$ at the contact points between the regions, which ultimately describes the entire spacetime through which the particles propagate.
@@
$dS=-2dr_{s}v_{s}f\left\{\frac{1}{1-v^{2}_{s}f^{2}}\right\}$ (36)
@@
is obtained, whose solution will be presented later (see Section 5), allowing us to obtain the expression for the quantum potential in region $II$. Thus, by integrating Eq. (36), as shown in the part $I$ of the Appendix, the phase function is obtained:
$S_{II}(x,x_{s}(t))=\frac{\alpha_{0}\beta_{0}}{2\sigma\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}}\Biggl\{\frac{\beta_{1}-\mu_{0}}{\sqrt{\mu_{0}}}\tan^{-1}{\Biggl[\frac{\tan{(x^{\prime}/2)}}{\sqrt{\mu_{0}}}\Biggr]}$ (38)
$-\frac{\beta_{1}-\mu_{1}}{\sqrt{\mu_{1}}}\tan^{-1}{\Biggl[\frac{\tan{(x^{\prime}/2)}}{\sqrt{\mu_{1}}}\Biggr]}\Biggr\}.$
Above $\beta_{0}=[2(\alpha_{1}-1)]/[(\alpha_{1}-1)^{2}-\alpha^{2}_{0}]$, $\beta_{1}=(\alpha_{1}+1)/(\alpha_{1}-1)$, $\beta_{2}=[(\alpha_{1}+1)^{2}-\alpha^{2}_{0}]/[(\alpha_{1}-1)^{2}-\alpha^{2}_{0}]$, $\beta_{3}=[\alpha_{1}^{2}-1+\alpha^{2}_{0}]/[(\alpha_{1}-1)^{2}-\alpha^{2}_{0}]$, $\mu_{0}=\beta_{3}-\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}$, $\mu_{1}=\beta_{3}+\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}$ and
@@
$\Psi_{II}(x,t)=\sum_{l=1}^{2}C_{II}^{l}\exp{\{iS_{II}^{l}(x,t)\}},$ (41)
where $E^{1}$ and $E^{2}$ in (41) are two distinct energy values used in the superposition to model the tunneling dynamics. However, it is important to note that physically acceptable solutions for Eq. (41), leading to a consistent phase function (38), are those where $\alpha_{1}\neq 0$ or $\alpha_{1}\neq 1$. Now, to be able to construct a general solution to the problem, it is necessary to impose continuity conditions that are consistent with the physics of the problem.
For this reason, it is essential to emphasize that the phases $S_{j}$ and the amplitudes $\sqrt{\rho_{j}}$ defined by expressions (28)-(29) are functions that depend only on $x^{\mu}$, and therefore transform as scalars $S_{j}^{\prime}(x^{\mu\prime})=S_{j}(x^{\mu})$ and $\rho_{j}^{\prime}(x^{\mu\prime})=\rho_{j}(x^{\mu})$. So, under a coordinate transformation $x\rightarrow x^{\mu}$, $\Psi_{j}$ maintains its form, since $S_{j}$ and $\sqrt{\rho_{j}}$ are invariant. Similarly, the phase $S_{II}$ defined by (38) is derived from the metric $\mathcal{G}{\mu\nu}$ of expression (33), resulting in a scalar function that evidently transforms as a scalar; moreover, it depends on $r_{s}$, which is invariant.
On the other hand, the covariant derivative [^3] Following the line of clarifications made previously in sections 2 and 3, we observe that in this work $()_{;\nu}=\nabla_{\nu}()=\partial_{\nu}()+\Gamma_{\mu\nu}^{\rho}()$ is used to denote the covariant derivative. $\Psi_{j;\mu}=\partial_{\mu}\Psi_{j}+\Gamma_{\mu\nu}^{\rho}\Psi_{j}$, must be continuous at the boundaries. This is easily verifiable by recalling that for $j=I,III$ the symbols $\Gamma_{\mu\nu}^{\rho}$ are non-null only for $\mu=\rho=1$, $\nu=0$, which leads to $\Gamma_{10}^{1}=\partial_{0}g_{00}/2=0$ because $g_{00}=-\{3Q\}^{2/3}$ is constant, making the covariant derivative reduce to $\Psi_{j;x}=\partial_{x}\Psi_{j}$. In region $II$, the non-null $\Gamma_{\mu\nu}^{\rho}$ symbols (when a/2<x<a/2-a/2<x<a/2) are $\Gamma_{00}^{1}=v_{s}^{2}f\partial_{1}f/2$, $\Gamma_{01}^{0}=\Gamma_{10}^{1}=-v_{s}\partial_{1}f/2$, and as already stated, they are all null at the boundary since $v_{s}=0$, so the covariant derivative again reduces to $\Psi_{j;x}=\partial_{x}\Psi_{j}$.
Then, imposing the continuity conditions of the solution and its derivatives, the expressions obtained from Eqs. (28) and (41) must be substituted in
$\Psi_{k}(x,t)\biggl|_{(x_{k},t_{k})}=\Psi_{k+I}(x,t)\biggl|_{(x_{k},t_{k})},$ (42)
$[\partial_{x}\Psi_{k}(x,t)]\biggl|_{(x_{k},t_{k})}=[\partial_{x}\Psi_{k+I}(x,t)]\biggl|_{(x_{k},t_{k})},$ (43)
with $x_{k}\in[-a/2,a/2]$, $t_{k}\in[t_{0},t_{1}]$ (where $t_{0}$ and $t_{1}$ mark the particle’s arrival at the boundaries x=a/2x=-a/2 (entry) and $x=a/2$ (exit), respectively) and $k=I,II,III$, we obtain the system of equations:
$\mathcal{C}_{I}^{1}\zeta_{1}+\mathcal{C}_{I}^{2}\zeta_{2}e^{i\Delta Et_{0}}$ $=$ $\zeta_{3}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{0}}],$
$\mathcal{C}_{I}^{1}\zeta_{1}^{\prime}+\mathcal{C}_{I}^{2}\zeta_{2}^{\prime}e^{i\Delta Et_{0}}$ $=$ $\zeta_{3}^{\prime}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{0}}],$ (44)
$\mathcal{C}_{III}^{1}\zeta_{5}+\mathcal{C}_{III}^{2}\zeta_{6}e^{i\Delta Et_{1}}$ $=$ $\zeta_{4}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{1}}],$
$\mathcal{C}_{III}^{1}\zeta_{5}^{\prime}+\mathcal{C}_{III}^{2}\zeta_{6}^{\prime}e^{i\Delta Et_{1}}$ $=$ $\zeta_{4}^{\prime}[\mathcal{C}_{II}^{1}+\mathcal{C}_{II}^{2}e^{i\Delta Et_{1}}].$
@@
$\mathcal{C}_{I}^{1}$ $=$ $\varpi_{4}e^{-i\Delta Et_{0}}[(e^{i\Delta Et_{0}}-1)\mathcal{C}_{II}^{1}+1],$
$\mathcal{C}_{I}^{2}$ $=$ $\varpi_{3}[(e^{i\Delta Et_{0}}-1)\mathcal{C}_{II}^{1}+1],$
$\mathcal{C}_{II}^{1}$ $=$ $-\frac{e^{i\Delta Et_{1}}}{e^{i\Delta Et_{1}}-1}\left\{\varpi_{2}\mathcal{C}_{III}^{1}-1\right\},$ (45)
$\mathcal{C}_{II}^{2}$ $=$ $\frac{1}{e^{i\Delta Et_{1}}-1}\left\{\varpi_{2}e^{i\Delta Et_{1}}\mathcal{C}_{III}^{1}-1\right\},$
$\mathcal{C}_{III}^{2}$ $=$ $\varpi_{1}e^{i\Delta Et_{1}}\mathcal{C}_{III}^{1}.$
@@
$\zeta_{l}$ $=$ $\sqrt{\rho_{I}^{l}(x,t)}\exp{\{iS_{I}^{l}(x,t)\}}\biggl|_{(-\frac{a}{2},t_{0})},$
$\zeta_{l}^{\prime}$ $=$ $\frac{k^{l}[(\mathcal{B}^{2}-\mathcal{A}^{2})\sin{\{2p_{\mu}x^{\mu}\}}+i2\mathcal{A}\mathcal{B}]}{2\sqrt{\rho_{I}^{l}(x,t)}}\exp{\{iS_{I}^{l}(x,t)\}}\biggl|_{(-\frac{a}{2},t_{0})},$
$\zeta_{l+2}$ $=$ $\exp{\{i[S_{II}(x,x_{s}(t))-V_{0}t]\}}\biggl|_{((-1)^{l+2}\frac{a}{2},t_{0})},$
$\zeta_{l+2}^{\prime}$ $=$ $\frac{i\alpha_{0}\beta_{0}\sec^{2}{(\frac{x^{\prime}}{2})}\tan{(2\sigma r_{s})}}{2\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}\sqrt{\cosh^{2}{(2\sigma r_{2})}-1}}$
$\times$ $\exp{\{i[S_{II}(x,x_{s}(t))-V_{0}t]\}}$
$\times$ $\biggl[\frac{\beta_{1}-\mu_{1}}{\mu_{1}+\tan^{2}{(\frac{x^{\prime}}{2})}}-\frac{\beta_{1}-\mu_{0}}{\mu_{0}+\tan^{2}{(\frac{x^{\prime}}{2})}}\biggr]\biggl|_{((-1)^{l+2}\frac{a}{2},t_{0})},$
$\zeta_{l+4}$ $=$ $\exp{\{iS_{III}^{l}(x,t)\}}\biggl|_{(\frac{a}{2},t_{0})},$
$\zeta_{l+4}^{\prime}$ $=$ $ik^{l}\exp{\{iS_{III}^{l}(x,t)\}}\biggl|_{(\frac{a}{2},t_{0})}.$
In the expressions above, values for $l=1,2$ are considered, along with the energy variation $\Delta E=E^{2}-E^{1}$. Additionally, the coefficients were rewritten as $C_{j}^{l}=\mathcal{C}_{j}^{l}$ if $j=I,II$ and $C_{j}^{l}\sqrt{\rho_{j}^{l}}=\mathcal{C}_{j}^{l}$ for $j=III$, again for all $l=1,2$.
@@
$\Psi_{I}(x,t)$ $=$ $\mathcal{C}_{I}^{1}\sqrt{\rho_{I}^{1}}\exp{\{iS_{I}^{1}(x,t)\}}+\mathcal{C}_{I}^{2}\sqrt{\rho_{I}^{2}}\exp{\{iS_{I}^{2}(x,t)\}},$ (47)
$\Psi_{II}(x,t)$ $=$ $[\mathcal{C}_{II}^{1}\exp{\{iE^{1}t\}}+\mathcal{C}_{II}^{2}\exp{\{iE^{2}t\}}]\exp{\{i[S_{II}(x,x_{s}(t))-V_{0}t]\}},$
$\Psi_{III}(x,t)$ $=$ $\mathcal{C}_{III}^{1}\exp{\{iS_{III}^{1}(x,t)\}}+\mathcal{C}_{III}^{2}\exp{\{iS_{III}^{2}(x,t)\}},$
@@ 5 General Expression for the Quantum Potential
For this purpose, it is necessary to recall that the metric describing spacetime in regions $I$ and $III$ implies that the quantum potential in these regions must be of the form $Q_{j}=\pm mc^{2}\{g_{00}\}^{3/2}/3=\pm\{g_{00}\}^{3/2}/3$, where again $j=I$ as $x\leq-a/2$ and $j=III$ as $x\geq a/2$. However, as mentioned in the previous section, the quantum potential in region II is determined by integrating Eq. (37), resulting in the expression
$Q_{II}(x,t)=\Biggl\{\ln{\Biggl|\frac{1+v_{s}f}{\sqrt{1-v^{2}_{s}f^{2}}}\Biggr|}+\frac{v^{2}_{s}f^{2}}{2}\Biggr\}.$ (49)
@@ 6 Momentum, Position, and Tunneling Time
In the previous sections, along with the theoretical foundations and the general solution, we derived the mathematical expressions for the quantum potential in each of the regions of interest in our problem. Even though the wave function and the quantum potential are fundamental quantities within the framework of Bohmian mechanics [36, 38, 47, 48], it is essential to determine the particle’s momentum and position. Therefore, this section is dedicated to calculating the mathematical expressions for these quantities, which ultimately allows us to determine the quantum tunneling time [1, 8, 9, 12, 16, 17, 18, 33].
In this regard, obtaining an expression for the linear momentum for the region II turns out to be a difficult task due to the form of the general solution. Nevertheless, by rewriting $\sqrt{\rho_{I}^{l}}$ in the form
$\sqrt{\rho_{I}^{l}}$ $=$ $\sqrt{\mathcal{A}^{2}\cos^{2}{\{p_{\mu}x^{\mu}/\hbar\}}+\mathcal{B}^{2}\sin^{2}{\{p_{\mu}x^{\mu}/\hbar\}}}$
$=$ $\sqrt{2AB+[A^{2}+B^{2}]\cos{\{2p_{\mu}x^{\mu}\}}}$
@@
$\sqrt{\rho_{I}^{l}}$ $=$ $\sqrt{2AB+[A^{2}+B^{2}]\cos{\{2p_{\mu}x^{\mu}\}}}$
$=$ $\sqrt{(A+B)^{2}}$
$=$ $\mathcal{A}.$
On the other hand, in the low-energy limit ($E\to 0$), the phase function $S_{I}^{l}$ simplifies to $S_{I}^{l}(x,t)=\tan^{-1}\left(\{\mathcal{B}\tan(p_{\mu}x^{\mu})/\mathcal{A}\}\right)\approx\mathcal{B}(k^{l}x-E^{l}t)/\mathcal{A}$, where $p_{\mu}x^{\mu}=k^{l}x-E^{l}t$ and the approximation holds for small arguments. When substituted into the superposition (29), this yields the emergent phase function $S_{I}(x,t)$, whose gradient governs the Bohmian velocity field in the low-energy regime.
To determine the linear momentum in region $I$, we choose the metric in this region to be of the Minkowski type, i.e., $Q_{I}=mc^{2}/3$, which implies that Eq. (20) reduces to Eq. (7) when the particles incident on the potential barrier have non-relativistic velocities ($v\ll 1$, in natural units $\hbar=c=1$), which is precisely our case of study. Then, substituting $S_{I}(x,t)$ into Eq. (7), we obtain
$P_{I}(x,t)$ $=$ $\frac{\mathcal{B}[(\mathcal{C}_{I}^{1})^{2}k^{1}+(\mathcal{C}_{I}^{2})^{2}k^{2}+\mathcal{C}_{I}^{1}\mathcal{C}_{I}^{2}(k^{1}+k^{2})\cos{\Xi_{I}}]}{\mathcal{A}[(\mathcal{C}_{I}^{1})^{2}+(\mathcal{C}_{I}^{2})^{2}+2\mathcal{C}_{I}^{1}\mathcal{C}_{I}^{2}\cos{\Xi_{I}}]}.$ (52)
@@
$P_{III}(x,t)$ $=$ $\frac{(\mathcal{C}_{III}^{1})^{2}k^{1}+(\mathcal{C}_{III}^{2})^{2}k^{2}+\mathcal{C}_{III}^{1}\mathcal{C}_{III}^{2}(k^{1}+k^{2})\cos{\Xi_{III}}}{(\mathcal{C}_{III}^{1})^{2}+(\mathcal{C}_{III}^{2})^{2}+2\mathcal{C}_{III}^{1}\mathcal{C}_{III}^{2}\cos{\Xi_{III}}}.$ (53)
@@
$P_{II}(x,t)$ $=$ $\frac{\alpha_{0}\beta_{0}\mbox{sech}(2\sigma r_{s})}{2\sqrt{(\beta_{3})^{2}-(\beta_{2})^{2}}}$
$\times$ $\Big\{\frac{1-\alpha_{0}^{2}+2\alpha_{1}\cosh{(2\sigma r_{s})}}{1+2\alpha_{1}\cosh{(2\sigma r_{s})}+\alpha_{1}^{2}\cosh^{2}{(2\sigma r_{s})}}$
$+$ $\frac{\alpha_{1}^{2}\cosh^{2}{(2\sigma r_{s})}}{1+2\alpha_{1}\cosh{(2\sigma r_{s})}+\alpha_{1}^{2}\cosh^{2}{(2\sigma r_{s})}}\Big\}$
$\times$ $\Biggl\{\frac{\beta_{1}-\mu_{0}}{\mu_{0}\cos^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}+\sin^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}}$
- $\frac{\beta_{1}-\mu_{1}}{\mu_{1}\cos^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}+\sin^{2}{[\frac{\mbox{sech}(2\sigma r_{s})}{2}]}}\Biggr\}.$